OpenStudy (anonymous):

Samantha is going on vacation for the summer and is trying to choose between two different plans. The first plan costs $450 for 3 days at a hotel and 2 days at an amusement park. The second plan offers 5 days at the same hotel and 3 days at the amusement park for $700. The cost of 1 day at the hotel and the amusement park is the same under both plans. How much does a 1 day trip to the amusement park cost?

OpenStudy (anonymous):

I'm really stuck. these are my options $50 $75 $150 $200

OpenStudy (keana):

Let h = cost of hotel per night and a = cost of amusement park for one day. Then we have two equations as follows: $450 = 3h + 2a $700 = 5h + 3a multiply first eqn by 5 and second by -3 $2250 = 15h + 10 a -$2100 = -15h - 9a Now add them: $150 = a The cost of one day at the amusement park is $150.

OpenStudy (anonymous):

Let x = hotel expense Let y = amusement park expense 3x + 2y = 450 5x + 3y = 700 -5(3x + 2y = 450) 3(5x + 3y = 700 -15x - 10y = -2250 15x + 9y = 2100 -y = -150 y = 150 The expense for 1 day at the amusement park is 150

OpenStudy (anonymous):

Thank you soooo much that that helped me understand it big time!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

can you guys help me on this one as well. I would really appreciate it! Solve 3x - 5y = 17 9x - 15y = -4 (-1, -4) (4, -1) No Solutions Infinitely Many Solutions

OpenStudy (anonymous):

3x - 5y = 17 9x - 15y = -4 -3(3x - 5y = 17) 9x - 15y = -4 -9x + 15y = -51 9x - 15y = -4 0 = -55 Therefore, no solution. THese are parallel lines.

OpenStudy (anonymous):

Okay that is what I thought it was as well so thank you for helping me make sure. :)

OpenStudy (anonymous):

You're welcome. You know when it's infinitely many solutions right?